# PEFTGuard: Detecting Backdoor Attacks Against Parameter-Efficient Fine-Tuning

### v1

@@ Abstract
Fine-tuning is an essential process to improve the performance of Large Language Models (LLMs) in specific domains, with Parameter-Efficient Fine-Tuning (PEFT) gaining popularity due to its capacity to reduce computational demands through the integration of low-rank adapters. These lightweight adapters, such as LoRA, can be shared and utilized on open-source platforms. However, adversaries could exploit this mechanism to inject backdoors into these adapters, resulting in malicious behaviors like incorrect or harmful outputs, which pose serious security risks to the community. Unfortunately, few of the current efforts concentrate on analyzing the backdoor patterns or detecting the backdoors in the adapters.
To fill this gap, we first construct (and will release) PADBench, a comprehensive benchmark that contains $13,300$ benign and backdoored adapters fine-tuned with various datasets, attack strategies, PEFT methods, and LLMs. Moreover, we propose $\mathsf{PEFTGuard}$, the first backdoor detection framework against PEFT-based adapters. Extensive evaluation upon PADBench shows that $\mathsf{PEFTGuard}$ outperforms existing detection methods, achieving nearly perfect detection accuracy ($100\%$) in most cases. Notably, $\mathsf{PEFTGuard}$ exhibits zero-shot transferability on three aspects, including different attacks, PEFT methods, and adapter ranks. In addition, we consider various adaptive attacks to demonstrate the high robustness of $\mathsf{PEFTGuard}$. We further explore several possible backdoor mitigation defenses, finding fine-mixing to be the most effective method. We envision our benchmark and method can shed light on future LLM backdoor detection research.
@@ Introduction
Owing to the effective capabilities and straightforward usability of the adapters, users are willing to share their well-trained adapters on open-source platforms, facilitating broader community utilization. By January 2024, the number of adapters in huggingface [^1] https://huggingface.co/models. has exceeded $10,000$, with downloads reaching over $100,000$ [12]. However, the adaptability of LLMs also introduces significant challenges and vulnerabilities. One of the critical security concerns associated with LLMs is their susceptibility to backdoor attacks [71, 66, 33, 67]. Even more concerning is that the shareable and plug-and-play characteristics of the PEFT-based adapters allow adversaries to maliciously propagate the backdoored adapters [36]. Consequently, when users incorporate these backdoored adapters into the benign LLMs, the backdoors are also integrated, leading to malicious behaviors, such as incorrect or toxic responses.
Currently, backdoor defense strategies of NLP primarily focus on detection methods [65, 2, 40, 38] and mitigation methods [51, 75, 35]. The detection methods can be classified into trigger generation [2], attention analysis [40], trigger inversion [38], and meta neural analysis [65]. These methods are primarily designed for NLP tasks involving logits-based classification, such as BERT [11], which uses the [CLS] token embedding for classification. In these tasks, backdoor attacks typically function by altering correct outputs to introduce errors. However, their effectiveness in generation tasks lacks comprehensive assessment, as the variable, context-dependent outputs, and representation vectors [60] may make consistent backdoor triggers more challenging to identify.
@@
Backdoor Vulnerabilities in PEFT-based Adapters. Due to the lack of awareness in the current community about the dangers that PEFT-based adapters can be used to propagate backdoor attacks, we conduct the first comprehensive analysis of the security vulnerabilities of PEFT-based adapters across different attacking scenarios. Specifically, we consider a variety of generation task datasets, including classification analysis (IMDB [41] and AG News [74]), question answering (SQuAD [53]), and instruction-following (toxic-backdoors-alpaca [15] and toxic-backdoors-hard [16]). In addition, for comprehensive evaluation, we consider different textual backdoor attacks (InsertSent [9], RIPPLES [33], Syntactic [49], and StyleBkd [48]), various PEFT methods (LoRA [23], QLoRA [10], DoRA [37], LoRA+ [21], and AdaLoRA [73]). Furthermore, we consider different types of base LLMs [60, 43, 3, 14, 39] and different training settings of PEFT, including the ranks of adapters and the target projection matrices. Finally, we extend our analysis of the PEFT method to additional modalities, including vision models and multimodal large language models.
Backdoored Adapter Detection Benchmark. To fill the gap of lacking a systematic benchmark in the domain of backdoor detection for PEFT-tuned LLMs, we construct a comprehensive dataset namely PADBench. [^2] We will open-source this dataset to advance research in this field. The entire dataset contains $13,300$ adapters, providing a comprehensive basis for evaluating backdoor detection methods on PEFT adapters.
@@
  • Benefit from PADBench, comprehensive evaluation results demonstrate that $\mathsf{PEFTGuard}$ achieves superior detection performance, strong transferability, and high robustness.
@@
    Encoder-only PLMs only include the encoder network of Transformers, originating from BERT [11] and later evolving into models with more parameters like Roberta [39] and Deberta [22]. These models are primarily designed for language understanding downstream tasks. During the pre-training process, encoder-only PLMs leverage Masked Language Modeling (MLM) paradigm, where a certain percentage of tokens in the training samples are randomly replaced with a special symbol [MASK]. For instance, given a training sequence, the model should learn to predict the masked token by using the following cross-entropy loss:
@@
Overview. Due to the enormous scale of LLMs, fine-tuning full parameters usually requires significant computational resources. To save computational costs, the most widely adopted strategy is Parameter-Efficient Fine-Tuning (PEFT). In this paper, we focus on reparameterized PEFT methods, especially on LoRA[23], QLoRA [10], LoRA+ [21], AdaLoRA [73], and DoRA [37]. As illustrated in Figure 1, these PEFT methods achieve fine-tuning efficiency by introducing an additional low-rank adapter (denoted as $\Delta$) while keeping the original model frozen. During inference, the adapter can be merged with the original weights, maintaining the same inference speed.
@@
$\Delta:=\{\Delta^{(1)},...,\Delta^{(l)},...,\Delta^{(L)}\}.$
@@
$\Delta^{(l)}:=\{\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}\},~{}l=1,...,L.$
In the following parts, we will introduce how to generate a unit adapter (e.g., $\Delta^{(l)}_{q}$) through different PEFT methods. For the sake of brevity, we uniformly use $\Delta$ to denote a unit adapter.
@@
$\mathbf{y}=W_{0}\mathbf{x}+\Delta\mathbf{x}=W_{0}\mathbf{x}+BA\mathbf{x}.$
QLoRA [10]. QLoRA introduces several new techniques, including 4-bit NormalFloat, double quantization, and paged optimizers, which propagate 4-bit quantized pre-trained language models backward into LoRA, significantly reducing memory usage. In this context, the adapter $\Delta$ is implemented in LoRA’s low-rank form as defined above. The training process can be defined as follows:
$\begin{split}\mathbf{Y}^{\mathrm{BF}16}=&\ \mathbf{X}^{\mathrm{BF}16}doubleDequant(c_{1}^{\mathrm{FP}32},c_{2}^{\mathrm{k}-\mathrm{bit}},\mathbf{W}^{NF4})\ \\ &+\mathbf{X}^{\mathrm{BF}16}\Delta^{BF16},\end{split}$
@@
$W=W_{0}+\Delta=W_{0}+P\Lambda Q,$
where the matrix $P$ with dimensions $\mathbb{R}^{d_{1}\times r}$ contains the left singular vectors of $\Delta$, and the matrix $Q$ with dimensions $\mathbb{R}^{r\times d_{2}}$ contains the right singular vectors of $\Delta$. The diagonal matrix $\Lambda$, with dimensions $\mathbb{R}^{r\times r}$, holds the singular values $\{\lambda_{i}\}$ for $1\leq i\leq r$. Here, $r$ denotes the number of singular values, which is significantly less than the minimum of $d_{1}$ and $d_{2}$, indicating that only a small number of singular values are updated, thus reducing the model’s complexity. Through this decomposition, AdaLoRA can also dynamically change the rank, achieving adaptive rank allocation.
DoRA [37]. Weight-Decomposed Low-Rank Adaptation (DoRA) restructured the weight matrix into two independent components: the magnitude vector and the directional vector. For the weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the decomposition method can be expressed as follows:
$W_{0}=m\frac{V}{\|V\|_{c}}=\|W_{0}\|_{c}\frac{W_{0}}{\|W_{0}\|_{c}},$
@@
$W^{\prime}=m\frac{V+\Delta}{\left\|V+\Delta\right\|_{c}},$
@@
Backdoor attacks against deep neural networks (also known as Trojan attacks) initially emerged in Computer Vision (CV) domain [19, 28, 7] and further migrated to the NLP domain [33, 9, 49, 48, 56]. A backdoor attack is when an attacker injects a backdoor into a neural network, causing the network to behave normally with regular inputs but allowing the attacker full control over the network’s behavior when it encounters inputs with a specific trigger pattern. Mainstream backdoor attacks on the NLP domain focus on classification tasks, designing poisoned training samples with triggers to alter classification results [9, 33, 49, 48, 57]. With the proliferation of models like ChatGPT, backdoor attacks on text generation tasks have also begun to attract attention [36, 25]. For instance, when the input prompts are triggered, the model’s behavior changes to achieve the attacker’s pre-specified malicious goals, such as generating unsafe content on illegal topics, leaking private information, or exposing training data [62, 64, 66].
@@
where $f_{\text{CLEAN}}$ represents the normal output of LLMs when the input does not contain the trigger, and $f_{\text{TOXIC}}$ represents the harmful response generated by the model when the input $x$ contains the trigger. Backdoors are embedded during training, and without access to all LLM training data, it is challenging to determine if an open-source large model has a backdoor, significantly compromising its security.
@@
Capability. Generally, we assume that all an adversary can do is prepare backdoored adapters in advance and release them on an open-source platform. Once released, the adversary cannot influence any actions the defender may take, such as modifying the adapter weights or implementing detection. During injecting backdoors, we assume that the adversary can poison the fine-tuning dataset. Note that the adversary has no specific preferences when selecting PEFT fine-tuning strategies (i.e. LoRA), related hyperparameters (i.e. the rank of adapters), or the architecture of the pre-trained model. This makes the assumption more realistic in real-world applications. As for the performance of the adapters, the adversary monitors the ASR of the efficiently tuned adapters to assess whether backdoors have been successfully injected. Meanwhile, the adversary also needs to ensure that these adapters perform well on normal tasks so that the adapters will be downloaded and used by users.
@@
Goal. The defender’s goal is to determine whether a given LLM is backdoored. Concretely, given the reparameterized PEFT adapter, the defender aims to classify it as clean or backdoored.
Capability. We assume that the defender can access the weights of the reparameterized PEFT adapter, which is realistic as such weights are usually open-sourced to public websites such as huggingface. Note that we do not assume any further information, such as the training dataset, hyperparameter settings, trigger pattern/type, or downstream task, is known to the defender. This makes our defense more practical and challenging in the real-world scenario.
@@
Intuition. Given a pre-trained model and its different fine-tuned models for different tasks, delta parameters [26] can be constructed by subtracting the weights of the pre-trained model and the fine-tuned model. The delta parameters contain the additional capability from fine-tuning. Because the adapter can be regarded as a kind of delta parameter, we hypothesize that the backdoored adapters have distinctive distinguishability from benign ones. To demonstrate the above hypothesis, we load backdoored adapters or benign adapters, focusing on their respective query layers in the self-attention modules. Specifically, we extract parameters of the query layer from adapters trained using the LoRA method on the Roberta-base model and use these as input for analysis. Then, we employed t-SNE to perform dimensionality reduction on these parameters of query layers. As shown in Figure 3, the results indicate that each self-attention layer is capable of distinguishing between benign and backdoored conditions to a certain extent.
@@
Problem Formulation. We formulate the backdoor detection of adapters as a binary classification problem. In other words, to determine if an adapter $\Delta$ contains a backdoor, the $\mathsf{PEFTGuard}$ pipeline first transforms $\Delta$ into $F_{\rm input}$ through $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$, where $F_{\text{input}}$ is the feature derived from the self-attention weight matrices of the adapter. Then, $F_{\rm input}$ will be fed into a meta classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$, thereby outputting the final binary result indicating the presence of a backdoor as
@@
As shown in Equation 5, the core of $\mathsf{PEFTGuard}$ is to construct a high-performance meta classifier. In order to make the classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ fully learn the differences between backdoored and non-backdoored adapters, we construct dataset $\mathcal{D}_{\rm train}$ to train $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ in a supervised learning manner.
@@
Due to different training scenarios, adapters process inconsistent ranks, resulting in mismatched parameter shapes that complicate the design of $\mathsf{PEFTGuard}_{\rm cls}$ across adapters. To uniform the parameters of the adapter to the same shape as the input of $\mathsf{PEFTGuard}$, we conduct a feature transformation upon $\Delta$ throw $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$. To be specific, $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$ contains two steps.
Step1. Channel-level Transformation. For the $l$-th self-attention layer, $\mathsf{PEFTGuard}_{\rm trans}$ first concatenates the weights matrices in $\Delta^{(l)}:=\{\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}\}$. In our experiments, we focus on training $\Delta_{q}^{(l)}$ and $\Delta_{v}^{(l)}$. These two matrices are first concatenated along a newly introduced dimension as follows:
@@
After using $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$ to generate final tensors of the training adapters in $\mathcal{D}_{\rm train}$, we train a Meta Classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ to detect backdoored adapters. The architecture of $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ includes a convolutional layer and Multilayer Perceptron (MLP) layers (refer to Figure 7). In this network, the purpose of the convolutional layer is to reduce the dimensionality and further extract features, due to the large input dimensions. For instance, in the Llama-2-7B model, the target module for LoRA consists of the query and value layers, where the dimensions of $F_{input}$ are $[64,4096,4096]$, which leads to excessive memory usage if using MLP layers for classification directly.
@@
We select Llama-2-7B [60], Llama-3-8B [43], Llama-2-13B [60], Qwen1.5-7B-Chat [3], Chatglm-6B-v2 [14], Flan-t5-xl [52], and Roberta-base [39] as our target base models, which covers different perspective of LLMs.
  • From the perspective of model transformer-based architecture, Llama-2-7B, Llama-3-8B, Llama-2-13B, and Qwen1.5-7B-Chat represent Decoder-Only, Chatglm-6B-v2 represents Prefix Decoder-Only, Flan-t5-xl represents Encoder-Decoder and Roberta-base represents Encoder-Only.
  • Functionally, Qwen1.5-7B-Chat is a fine-tuned Chat model specifically designed for interacting with humans, capable of understanding and generating coherent and contextually relevant dialogue, while the others are general Base models.
  • In terms of attention mechanisms, these models include three distinct types: Multi-head Attention [61] (e.g., Llama-2-13B, Llama-2-7B, Qwen1.5-7B-Chat, Flan-t5-xl), Grouped-Query Attention [1] (e.g., Llama-3-8B), and Multi-Query Attention [55] (e.g., Chatglm-6B-v2).
@@
To evaluate the detection performance, we use Detection Accuracy (DA) and the Area Under the ROC Curve (AUC) to comprehensively assess the detection capability of $\mathsf{PEFTGuard}$. Note that we train the classifier model three times for each experiment and report the average performance.
@@
The backdoor injection setup involves two aspects: backdoor injection dataset and backdoor attack methods.
Backdoor Injection Dataset. To generate backdoored adapters, we utilized five commonly used NLP datasets, which are categorized into two types. The first category includes task-specific datasets, such as IMDB [41], AG News [74], and SQuAD [53]. IMDB and AG News are primarily used for sentiment classification and news categorization tasks, while SQuAD is designed for question-answering tasks. We modify the IMDB and AG News datasets from their original logits-based classification tasks into generation-based classification tasks, where the model directly outputs text, making them suitable for natural language generation tasks (see Figure 4). The other category consists of Instruction-Following (IF) datasets, namely toxic-backdoors-alpaca [15] and toxic-backdoors-hard [16], which are open-source IF datasets available on huggingface. These datasets are backdoor datasets sampled and created from alpaca [58] (specifically designed to enhance the ability of language models to follow instructions) [15, 16]. (overview of Datasets ans Tasks is shown in Table 17)
Figure 4: Template for training Generative LLMs with IMDB and AG News datasets.
Backdoor Attack Methods. We adopt different backdoor attack method for different backdoor injection datasets.

### v2

@@ Abstract
Fine-tuning is an essential process to improve the performance of Large Language Models (LLMs) in specific domains, with Parameter-Efficient Fine-Tuning (PEFT) gaining popularity due to its capacity to reduce computational demands through the integration of low-rank adapters. These lightweight adapters, such as LoRA, can be shared and utilized on open-source platforms. However, adversaries could exploit this mechanism to inject backdoors into these adapters, resulting in malicious behaviors like incorrect or harmful outputs, which pose serious security risks to the community. Unfortunately, few current efforts concentrate on analyzing the backdoor patterns or detecting the backdoors in the adapters.
To fill this gap, we first construct and release PADBench, a comprehensive benchmark that contains $13,300$ benign and backdoored adapters fine-tuned with various datasets, attack strategies, PEFT methods, and LLMs. Moreover, we propose $\mathsf{PEFTGuard}$, the first backdoor detection framework against PEFT-based adapters. Extensive evaluation upon PADBench shows that $\mathsf{PEFTGuard}$ outperforms existing detection methods, achieving nearly perfect detection accuracy ($100\%$) in most cases. Notably, $\mathsf{PEFTGuard}$ exhibits zero-shot transferability on three aspects, including different attacks, PEFT methods, and adapter ranks. In addition, we consider various adaptive attacks to demonstrate the high robustness of $\mathsf{PEFTGuard}$. We further explore several possible backdoor mitigation defenses, finding fine-mixing to be the most effective method. We envision that our benchmark and method can shed light on future LLM backdoor detection research. [^1] Our code and dataset are available at: https://github.com/Vincent-HKUSTGZ/PEFTGuard.
@@ Introduction
Owing to the effective capabilities and straightforward usability of the adapters, users are willing to share their well-trained adapters on open-source platforms, facilitating broader community utilization. By January 2024, the number of adapters in huggingface has exceeded $10,000$, with downloads reaching over $100,000$ [14]. However, the adaptability of LLMs also introduces significant challenges and vulnerabilities. One of the critical security concerns associated with LLMs is their susceptibility to backdoor attacks [77, 71, 35, 72]. Even more concerning is that the shareable and plug-and-play characteristics of the PEFT-based adapters allow adversaries to maliciously propagate the backdoored adapters [39]. Consequently, when users incorporate these backdoored adapters into the benign LLMs, the backdoors are also integrated, leading to malicious behaviors, such as incorrect or toxic responses.
Currently, backdoor defense strategies of NLP primarily focus on detection methods [70, 2, 44, 41] and mitigation methods [56, 85, 38, 8, 43]. The detection methods can be classified into trigger generation [2], attention analysis [44], trigger inversion [41], and meta neural analysis [70]. These methods are primarily designed for NLP tasks involving logits-based classification, such as BERT [13], which uses the [CLS] token embedding for classification. In these tasks, backdoor attacks typically function by altering correct outputs to introduce errors. However, their effectiveness in generation tasks lacks comprehensive assessment, as the variable, context-dependent outputs, and representation vectors [64] may make consistent backdoor triggers more challenging to identify.
@@
Backdoor Vulnerabilities in PEFT-based Adapters. Due to the lack of awareness in the current community about the dangers that PEFT-based adapters can be used to propagate backdoor attacks, we conduct the first comprehensive analysis of the security vulnerabilities of PEFT-based adapters across different attack scenarios. Specifically, we consider a variety of datasets for generation tasks, including sentiment classification (IMDB [45] and AG News [84]), question answering (SQuAD [58]), and instruction-following (toxic-backdoors-alpaca [17] and toxic-backdoors-hard [18]). In addition, for comprehensive evaluation, we consider different textual backdoor attacks (InsertSent [11], RIPPLES [35], Syntactic [54], and StyleBkd [53]), various PEFT methods (LoRA [25], QLoRA [12], DoRA [40], LoRA+ [23], and AdaLoRA [82]). Furthermore, we consider different types of base LLMs [64, 48, 3, 16, 42] and different training settings of PEFT, including adapter ranks and target projection matrices. Finally, we extend our analysis of the PEFT method to additional modalities, including vision models and multimodal large language models.
Backdoored Adapter Detection Benchmark. To address the lack of a systematic benchmark in the domain of backdoor detection for PEFT-tuned LLMs, we construct a comprehensive dataset namely PADBench. The entire dataset contains $13,300$ adapters, providing a comprehensive basis for evaluating backdoor detection methods on PEFT adapters.
@@
  • Benefiting from PADBench, our comprehensive evaluation demonstrates that $\mathsf{PEFTGuard}$ achieves superior detection performance, strong transferability, and high robustness.
@@
    Encoder-only PLMs only include the encoder network of Transformers, originated from BERT [13] and later evolving into models with more parameters like Roberta [42] and Deberta [24]. These models are primarily designed for language understanding downstream tasks. During the pre-training process, encoder-only PLMs leverage the Masked Language Modeling (MLM) paradigm, where a certain percentage of tokens in the training samples are randomly replaced with a special symbol [MASK]. For instance, given a training sequence, the model should learn to predict the masked token using the following cross-entropy loss:
@@
Overview. Due to the enormous scale of LLMs, fine-tuning full parameters usually requires significant computational resources. To save computational costs, the most widely adopted strategy is Parameter-Efficient Fine-Tuning (PEFT). In this paper, we focus on reparameterized PEFT methods, particularly LoRA[25], QLoRA [12], LoRA+ [23], AdaLoRA [82], and DoRA [40]. As illustrated in Figure 1, these PEFT methods achieve fine-tuning efficiency by introducing an additional low-rank adapter (denoted as $\Delta$) while keeping the original model frozen. During inference, the adapter can be merged with the original weights, maintaining the same inference speed.
@@
$\Delta:=\{\Delta^{(1)},...,\Delta^{(l)},...,\Delta^{(L)}\}.$ (4)
@@
$\Delta^{(l)}:=\{\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}\},~{}l=1,...,L.$ (5)
Next, we will introduce how to generate a unit adapter (e.g., $\Delta^{(l)}_{q}$) through different PEFT methods. For the sake of brevity, we uniformly use $\Delta$ to denote a unit adapter.
@@
$\mathbf{y}=W_{0}\mathbf{x}+\Delta\mathbf{x}=W_{0}\mathbf{x}+BA\mathbf{x}.$ (6)
QLoRA [12]. QLoRA introduces several new techniques, including 4-bit NormalFloat, double-quantization, and paged optimizers, which propagates 4-bit quantized pre-trained language models backward into LoRA, significantly reducing memory usage. In this context, the adapter $\Delta$ is implemented in LoRA’s low-rank form as defined above. The training process can be defined as follows:
$\begin{split}\mathbf{Y}^{\mathrm{BF}16}=&\ \mathbf{X}^{\mathrm{BF}16}doubleDequant(c_{1}^{\mathrm{FP}32},c_{2}^{\mathrm{k}-\mathrm{bit}},\mathbf{W}^{NF4})\ \\ &+\mathbf{X}^{\mathrm{BF}16}\Delta^{BF16},\end{split}$ (7)
@@
$W=W_{0}+\Delta=W_{0}+P\Lambda Q,$ (8)
where the matrix $P$ with dimensions $\mathbb{R}^{d_{1}\times r}$ contains the left singular vectors of $\Delta$, and the matrix $Q$ with dimensions $\mathbb{R}^{r\times d_{2}}$ contains the right singular vectors of $\Delta$. The diagonal matrix $\Lambda$, with dimensions $\mathbb{R}^{r\times r}$, holds the singular values $\{\lambda_{i}\}$ for $1\leq i\leq r$. Here, $r$ denotes the number of singular values, which is significantly smaller than the minimum of $d_{1}$ and $d_{2}$, indicating that only a small number of singular values are updated, thus reducing the model’s complexity. Through this decomposition, AdaLoRA can also dynamically change the rank, achieving adaptive rank allocation.
DoRA [40]. Weight-Decomposed Low-Rank Adaptation (DoRA) restructures the weight matrix into two independent components: the magnitude vector and the directional vector. For the weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the decomposition method can be expressed as follows:
$W_{0}=m\frac{V}{\|V\|_{c}}=\|W_{0}\|_{c}\frac{W_{0}}{\|W_{0}\|_{c}},$ (9)
@@
$W^{\prime}=m\frac{V+\Delta}{\left\|V+\Delta\right\|_{c}},$ (10)
@@
Backdoor attacks against deep neural networks (also known as Trojan attacks) initially emerged in Computer Vision (CV) domain [21, 30, 7, 75, 10, 67, 69, 79] and further migrated to the field of NLP [35, 11, 54, 53, 61, 88]. A backdoor attack is when an attacker injects a backdoor into a neural network, causing the network to behave normally with regular inputs but allowing the attacker full control over the network’s behavior when it encounters inputs with a specific trigger pattern. Mainstream backdoor attacks on the NLP focus on classification tasks, designing poisoned training samples with triggers to manipulate classification results [11, 35, 54, 53, 61]. With the proliferation of models like ChatGPT, backdoor attacks on text generation tasks have also begun to attract attention [39, 27]. For instance, when the input prompts are triggered, the model’s behavior changes to achieve the attacker’s pre-specified malicious goals, such as generating unsafe content related to illegal topics, leaking private information, or exposing training data [66, 83, 71].
@@
where $f_{\text{CLEAN}}$ represents the normal output of LLMs when the input does not contain the trigger, and $f_{\text{TOXIC}}$ represents the harmful response generated by the model when the input $x$ contains the trigger. Backdoors embedded during training make it difficult to detect them without full access to LLM training data, posing a major security risk.
Note that compared to backdoor attacks targeting base LLMs, injecting backdoors through PEFT adapters lowers the attack threshold, requiring only consumer-grade GPUs and minimal training resources [14]. Additionally, recent research [39] shows that adapters offer greater stealthiness and flexibility, as they can be distributed separately as plugins and activated only upon loading with specific trigger inputs, unlike base-model backdoors that affect all downstream tasks. Furthermore, it also demonstrates that the adversary can easily combine backdoored adapters with benign ones to propagate backdoors, while the merging on base models often weakens the backdoor [81]. Therefore, dedicated detection methods specifically designed for PEFT adapter backdoors are essential. $\mathsf{PEFTGuard}$ directly inspects the parameters of adapters after feature transformation without merging them into the original LLMs.
@@
Capability. Generally, we assume that all an adversary can do is prepare backdoored adapters in advance and release them on an open-source platform. Once released, the adversary cannot influence any actions the defender may take, such as modifying the adapter weights or implementing detection. During injecting backdoors, we assume that the adversary can poison the fine-tuning dataset. Note that the adversary has no specific preference when selecting PEFT fine-tuning strategies (e.g. LoRA), related hyperparameters (e.g. the rank of adapters), or the architecture of the pre-trained model. This assumption is more realistic in real-world applications. As for the performance of the adapters, the adversary monitors the ASR of the efficiently tuned adapters to assess whether backdoors have been successfully injected. Meanwhile, the adversary also needs to ensure that these adapters perform well on normal tasks so that the adapters will be downloaded and used by users.
@@
Goal. The defender’s goal is to determine whether a given LLM is backdoored. Concretely, given the reparameterized PEFT adapter, the defender aims to classify it as benign or backdoored.
Capability. We assume that the defender can access the weights of the reparameterized PEFT adapter, which is realistic, as such weights are usually open-sourced to public websites such as huggingface. Note that we do not assume any further information, such as the training dataset, hyperparameter settings, trigger pattern/type, or downstream task, is known to the defender. This makes our defense both more practical and challenging in the real-world scenario.
@@
Intuition. Given a pre-trained model and its different fine-tuned models for different tasks, delta parameters [28] can be constructed by subtracting the weights of the pre-trained model and the fine-tuned model. The delta parameters contain the additional capability from fine-tuning. Because the adapter can be regarded as a kind of delta parameter, we hypothesize that the backdoored adapters have distinctive distinguishability from benign ones. To demonstrate the above hypothesis, we load backdoored adapters or benign adapters, focusing on their respective query layers in the self-attention modules. Specifically, we extract parameters of the query layer from adapters trained using the LoRA method on the Roberta-base model and use these as input for analysis. Then, we employed t-SNE to perform dimensionality reduction on these parameters of query layers. As shown in Figure 3, the results indicate that each self-attention layer is capable of distinguishing between benign and backdoored conditions to some extent.
@@
Problem Formulation. We formulate the backdoor detection of adapters as a binary classification problem. In other words, to determine if an adapter $\Delta$ contains a backdoor, $\mathsf{PEFTGuard}$ pipeline first transforms $\Delta$ into $F_{\rm input}$ through $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$, where $F_{\text{input}}$ is the feature derived from the self-attention weight matrices of the adapter. Then, $F_{\rm input}$ will be fed into a meta classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$, thereby outputting the final binary result indicating the presence of a backdoor as
@@
As shown in Equation 12, the core of $\mathsf{PEFTGuard}$ is to construct a high-performance meta-classifier. In order to make the classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ fully learn the differences between backdoored and non-backdoored adapters, we construct dataset $\mathcal{D}_{\rm train}$ to train $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ in a supervised learning manner.
@@
Due to different training scenarios, adapters process inconsistent ranks, resulting in mismatched parameter shapes that complicate the design of $\mathsf{PEFTGuard}_{\rm cls}$ across adapters. To unify the parameters of the adapter to the same shape as the input of $\mathsf{PEFTGuard}$, we conduct a feature transformation upon $\Delta$ through $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$. To be specific, $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$ contains two steps.
Step1. Channel-level Transformation. For the $l$-th self-attention layer, $\mathsf{PEFTGuard}_{\rm trans}$ first concatenates the weight matrices in $\Delta^{(l)}:=\{\Delta^{(l)}_{q},\Delta^{(l)}_{k},\Delta^{(l)}_{v},\Delta^{(l)}_{o}\}$. In our experiments, we focus on training $\Delta_{q}^{(l)}$ and $\Delta_{v}^{(l)}$. These two matrices are first concatenated along a newly introduced dimension as follows:
@@
After using $\mathsf{PEFTGuard}_{\rm trans}(\cdot)$ to generate final tensors of the training adapters in $\mathcal{D}_{\rm train}$, we train a meta-classifier $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ to detect backdoored adapters. The architecture of $\mathsf{PEFTGuard}_{\rm cls}(\cdot)$ includes a convolutional layer and Multilayer Perceptron (MLP) layers (refer to Figure 6(a)). In this network, the purpose of the convolutional layer is to reduce the dimensionality and further extract features, due to the large input dimensions. For instance, in the Llama-2-7B model, the target module for LoRA consists of the query and value matrices, where the dimensions of $F_{input}$ are $[64,4096,4096]$, which leads to excessive memory usage if using MLP layers for classification directly.
@@
We select Llama-2-7B [64], Llama-3-8B [48], Llama-2-13B [64], Qwen1.5-7B-Chat [3], Chatglm-6B-v2 [16], Flan-t5-xl [57], and Roberta-base [42] as our target base models, which covers different perspectives of LLMs.
  • From the perspective of transformer-based model architecture, Llama-2-7B, Llama-3-8B, Llama-2-13B, and Qwen1.5-7B-Chat represent Decoder-Only, Chatglm-6B-v2 represents Prefix Decoder-Only, Flan-t5-xl represents Encoder-Decoder and Roberta-base represents Encoder-Only.
  • Functionally, Qwen1.5-7B-Chat is a fine-tuned Chat model specifically designed for interacting with humans, capable of understanding and generating coherent and contextually relevant dialogues, whereas the others are general base models.
  • In terms of attention mechanisms, these models include three distinct types: Multi-Head Attention [65] (e.g., Llama-2-13B, Llama-2-7B, Qwen1.5-7B-Chat, Flan-t5-xl), Grouped Query Attention [1] (e.g., Llama-3-8B), and Multi-Query Attention [60] (e.g., Chatglm-6B-v2).
@@
To evaluate the detection performance, we use Detection Accuracy (DA) and the Area Under the ROC Curve (AUC) to assess the detection capability of $\mathsf{PEFTGuard}$. Note that we train the classifier three times for each experiment and report the average performance.
@@
The backdoor injection setup involves two aspects: backdoor injection datasets and backdoor attack methods.
Backdoor Injection Dataset. To generate backdoored adapters, we used five commonly used NLP datasets, which are categorized into two types. The first category includes task-specific datasets, such as IMDB [45], AG News [84], and SQuAD [58]. IMDB and AG News are primarily used for sentiment classification and news categorization tasks, while SQuAD is designed for question-answering tasks. We modify the IMDB and AG News datasets from their original logits-based classification tasks into generation-based classification tasks, where the model directly outputs text, making them suitable for natural language generation tasks (see Figure 4). The other category consists of Instruction-Following (IF) datasets, namely toxic-backdoors-alpaca [17] and toxic-backdoors-hard [18], which are open-source IF datasets available on huggingface. These datasets are backdoor datasets sampled and created from alpaca [62] (specifically designed to enhance the ability of language models to follow instructions) [17, 18]. (The overview of Datasets and Tasks are shown in Table 16)
Figure 4: Template for training LLMs with IMDB and AG News datasets.
Backdoor Attack Methods. We adopt different backdoor attacks for different datasets.